Please can you help me to find the general solution in radians of the following equation sin+2x++-+1+=+cos+2x
I think I need to use a trig identity but I'm not sure which one or how?
Thank you

Answer by venugopalramana(3286)

sin 2x - 1 = cos 2x
sin 2x - cos 2x = 1...divide throughout with square root of sum of squares of coefficients of sin(2x) and cos(2x) ...we get sq.rt(1^1+1^2)=sq.rt(2)..dividing with sq.rt(2)
if we put sin(1/sqrt(2))=pi/4 = cos(pi/4)..we get
cos(pi/4)sin(2x)+sin(pi/4)cos(2x)=1
sin(pi/4+2x)=1=sin(pi/2)
general solution is
(pi/4)+2x=n*(pi)+{(-1)^n}*(pi/2)
x=(1/2){n*(pi)-(pi/4)+((-1)^n)*(pi/2)}